1 Front end side

Three racks to check. All NIM crate have to powered on. Check single rates on the scaler for the total number of triggers, it should be of the order 1.8 KHz.

Rack  Crate  On ?
T1 rate  Hz  OK?
RR1 NIM Ampli Sum Left
RR1  NIM Discriminators
RR1  NIM Trigger Logic
RR2 NIM Ampli Sum Veto
RR2  NIM Trigger Logic
RR3 NIM Ampli Sum Left
RR3  NIM Discriminators
RR3  NIM Trigger Logic

2 DAQ side

Rack spectrometer side Fastbus crate.

All cables plugged to patch panel. Racks other side Camac discriminators VME crate

Rack  Crate  On ?
RR1  CAMAC Left
RR2  CAMAC Veto
RR2 CAMAC F1 VME TDC
RR3  CAMAC Right

3 Things to be checked if something look wrong

3.1 Thresholds

Discriminator  Module  Value Read out value
Veto  CAMAC Left  ?  ?
Sums  NIM discriminator  50mV  50mV
Neutron  Nim Discriminator  20mV
TDC  CAMAC Discrimintor 200mV  2V